# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v1

@@ C.1 Weighted majority vote
The idea to combine conformal prediction sets by weighted majority vote was introduced by Gasparin and Ramdas [17] as an extension of the majority vote method first proposed by Cherubin [23]. At first consideration, weighted majority vote appears to differ from our method: weighted majority vote performs weighted aggregation of prediction sets, while our method performs weighted aggregation of the p-values associated with prediction sets. Despite this distinction, Appendix B of Gasparin and Ramdas [17] shows that these two methods are, in fact, dual to each other. Thus, under data-independent weights, our method is essentially the same as theirs. Nevertheless, our p-value formulation enables two key extensions that go beyond weighted majority vote.
@@ E.2 Mixture-of-Experts model
E.3 Expert feature assignment for UCI experiments
@@
SugarAndAlcohol residual_sugar, density, alcohol
SulfurAndSalinity chlorides, free_sulfur_dioxide, total_sulfur_dioxide, sulphates
@@
RaceEthnicity racepctblack, racePctWhite, racePctAsian, racePctHisp, PctForeignBorn, PctSpeakEnglOnly, PctNotSpeakEnglWell, PctBornSameState, PctSameHouse85, PctSameCity85
IncomePoverty medIncome, medFamInc, perCapInc, whitePerCap, blackPerCap, indianPerCap, AsianPerCap, HispPerCap, NumUnderPov, PctPopUnderPov
EmploymentIndustry pctWWage, pctWFarmSelf, pctWInvInc, pctWSocSec, pctWPubAsst, pctWRetire, PctUnemployed, PctEmploy, PctEmplManu, PctEmplProfServ
OccupationEducation PctOccupManu, PctOccupMgmtProf, PctWorkMomYoungKids, PctWorkMom, PctUsePubTrans, PctLess9thGrade, PctNotHSGrad, PctBSorMore, MalePctDivorce, MalePctNevMarr
@@
HousingQualityCosts PctPersDenseHous, PctVacantBoarded, PctVacMore6Mos, MedYrHousBuilt, PctWOFullPlumb, OwnOccLowQuart, OwnOccMedVal, OwnOccHiQuart, RentLowQ, RentMedian
HousingCostsHomelessness RentHighQ, MedRent, MedRentPctHousInc, MedOwnCostPctInc, MedOwnCostPctIncNoMtg, NumInShelters, NumStreet, NumIlleg, PctIlleg, LemasPctOfficDrugUn
@@
[…]

### v2

@@ C.1 Weighted majority vote
The idea to combine conformal prediction sets by weighted majority vote was introduced by Gasparin and Ramdas [17] as an extension of the majority vote method first proposed by Cherubin [23]. At first consideration, weighted majority vote appears to differ from our method: weighted majority vote performs weighted aggregation of prediction sets, while our method performs weighted aggregation of the p-values associated with prediction sets. Despite this distinction, Appendix B of Gasparin and Ramdas [17] observes that these two methods are, in fact, dual to each other under data-independent weights, the setting considered in their work. Nevertheless, our p-value formulation enables two key extensions that go beyond weighted majority vote.
@@ E.2 Mixture-of-experts model
E.3 Computing prediction sets
To compute prediction sets, we leverage what we know about the structure of the p-value function. Observe that for a fixed input $x$, the function $\hat{p}_{k}(x,y)$ (2) is piecewise constant with discontinuities only at the values of $y$ that solve $s_{k}(x,y)=s_{k}(X_{i},Y_{i})$. (For an absolute residual score function $s_{k}(x,y)=|y-\hat{\mu}_{k}(x)|$, these values are equal to $\hat{mu}_{k}(x)\pm s(X_{i},Y_{i})$ for $i\in S_{k}$.) For each test point $x$, $\hat{p}$ is therefore a step function of $y$ with finite discontinuities, and the weighted average p-value function is likewise a step function of $y$, with its discontinuities as a union of the $K$ separate sets of discontinuities.
We use the portion library [39] to represent and manipulate our prediction sets; this also allows us to compute their Lebesgue measures without any discretization.
E.4 Expert feature assignment for UCI experiments
@@
Sugar/alcohol residual_sugar, density, alcohol
Sulfur/salinity chlorides, free_sulfur_dioxide, total_sulfur_dioxide, sulphates
@@
Race/ethnicity racepctblack, racePctWhite, racePctAsian, racePctHisp, PctForeignBorn, PctSpeakEnglOnly, PctNotSpeakEnglWell, PctBornSameState, PctSameHouse85, PctSameCity85
Income/poverty medIncome, medFamInc, perCapInc, whitePerCap, blackPerCap, indianPerCap, AsianPerCap, HispPerCap, NumUnderPov, PctPopUnderPov
Employment/industry pctWWage, pctWFarmSelf, pctWInvInc, pctWSocSec, pctWPubAsst, pctWRetire, PctUnemployed, PctEmploy, PctEmplManu, PctEmplProfServ
Occupation/education PctOccupManu, PctOccupMgmtProf, PctWorkMomYoungKids, PctWorkMom, PctUsePubTrans, PctLess9thGrade, PctNotHSGrad, PctBSorMore, MalePctDivorce, MalePctNevMarr
@@
Housing quality PctPersDenseHous, PctVacantBoarded, PctVacMore6Mos, MedYrHousBuilt, PctWOFullPlumb, OwnOccLowQuart, OwnOccMedVal, OwnOccHiQuart, RentLowQ, RentMedian
Homelessness RentHighQ, MedRent, MedRentPctHousInc, MedOwnCostPctInc, MedOwnCostPctIncNoMtg, NumInShelters, NumStreet, NumIlleg, PctIlleg, LemasPctOfficDrugUn
@@
Atomic mass mean_atomic_mass, wtd_mean_atomic_mass, gmean_atomic_mass, wtd_gmean_atomic_mass, entropy_atomic_mass, wtd_entropy_atomic_mass, range_atomic_mass, wtd_range_atomic_mass, std_atomic_mass, wtd_std_atomic_mass, number_of_element
Atomic radius mean_atomic_radius, wtd_mean_atomic_radius, gmean_atomic_radius, wtd_gmean_atomic_radius, entropy_atomic_radius, wtd_entropy_atomic_radius, range_atomic_radius, wtd_range_atomic_radius, std_atomic_radius, wtd_std_atomic_radius
Density mean_Density, wtd_mean_Density, gmean_Density, wtd_gmean_Density, entropy_Density, wtd_entropy_Density, range_Density, wtd_range_Density, std_Density, wtd_std_Density
Electron affinity mean_ElectronAffinity, wtd_mean_ElectronAffinity, gmean_ElectronAffinity, wtd_gmean_ElectronAffinity, entropy_ElectronAffinity, wtd_entropy_ElectronAffinity, range_ElectronAffinity, wtd_range_ElectronAffinity, std_ElectronAffinity, wtd_std_ElectronAffinity
FIE mean_fie, wtd_mean_fie, gmean_fie, wtd_gmean_fie, entropy_fie, wtd_entropy_fie, range_fie, wtd_range_fie, std_fie, wtd_std_fie
Fusion heat mean_FusionHeat, wtd_mean_FusionHeat, gmean_FusionHeat, wtd_gmean_FusionHeat, entropy_FusionHeat, wtd_entropy_FusionHeat, range_FusionHeat, wtd_range_FusionHeat, std_FusionHeat, wtd_std_FusionHeat
Thermal conductivity mean_ThermalConductivity, wtd_mean_ThermalConductivity, gmean_ThermalConductivity, wtd_gmean_ThermalConductivity, entropy_ThermalConductivity, wtd_entropy_ThermalConductivity, range_ThermalConductivity, wtd_range_ThermalConductivity, std_ThermalConductivity, wtd_std_ThermalConductivity
Valence mean_Valence, wtd_mean_Valence, gmean_Valence, wtd_gmean_Valence, entropy_Valence, wtd_entropy_Valence, range_Valence, wtd_range_Valence, std_Valence, wtd_std_Valence